# IGEV++: Iterative Multi-range Geometry Encoding Volumes for Stereo Matching

### v1

@@ III-C ConvGRU-based Update Operator
We input the geometry features of MGEV into the ConvGRU-based update operator to iteratively update the disparity field. We apply $soft\;argmin$ to $\mathbf{G}^{s}$, $\mathbf{G}^{m}$, and $\mathbf{G}^{l}$ to regress a set of initial disparity maps $\mathbf{d}_{0}^{s}$, $\mathbf{d}_{0}^{m}$, and $\mathbf{d}_{0}^{l}$,
$\mathbf{d}_{0}^{s}=$ $\;\sum\limits_{d=0}^{D^{s}-1}d\times Softmax(\mathbf{G}^{s}(d)),$ (5)
$\mathbf{d}_{0}^{m}=$ $\;\sum\limits_{d=0}^{D^{m}/2-1}2d\times Softmax(\mathbf{G}^{m}(d)),$
$\mathbf{d}_{0}^{l}=$ $\;\sum\limits_{d=0}^{D^{l}/4-1}4d\times Softmax(\mathbf{G}^{l}(d)).$
To speed up convergence, we take the initial disparity map $\mathbf{d}^{s}_{0}$ as the starting point for ConvGRU-based iterations. In each GRU-based iteration, we utilize the previous disparity map to index a set of geometry features $\mathbf{f}_{G}^{s}$, $\mathbf{f}_{G}^{m}$, and $\mathbf{f}_{G}^{l}$ from $\mathbf{G}^{s}$, $\mathbf{G}^{m}$, and $\mathbf{G}^{l}$. We further propose a Selective Geometry Feature Fusion module (SGFF) to effectively integrate geometry features from different ranges and granularities of MGEV.
@@
$\mathbf{s}_{i}=$ $\;\sigma(\text{Conv}(\text{Concat}\{\mathbf{f}_{l,4},\mathbf{f}_{d}\})),i=s,m,l.$
@@
In each GRU-based iteration, the fused geometry feature $\mathbf{f}_{G}$ and the previous disparity map $\mathbf{d}_{k-1}$ are passed through two encoder layers and then concatenated with $\mathbf{d}_{k-1}$ to form $x_{k}$. Then we use ConvGRUs to update the hidden state $h_{k-1}$ as RAFT-Stereo [3],
@@
where $W_{z}$, $W_{r}$, and $W_{h}$ are the parameters of the network, $c_{z}$, $c_{r}$, $c_{h}$ are context features generated from the context network. The number of channels for both the hidden states and the context features is 128. The $\text{Encoder}_{g}$ and $\text{Encoder}_{d}$ consist of two convolutional layers respectively. Based on the hidden state $h_{k}$, we decode a residual disparity $\triangle\mathbf{d}_{k}$ through two convolutional layers, then we update the current disparity map,
@@
Figure 6: Visual comparison with RAFT-Stereo [3] using few iterations for inference, with iteration numbers of 1, 2, and 3.
@@ IV-A Datasets and Evaluation Metrics
Middlebury 2014 [18] is an indoor dataset, which provides 15 training pairs and 15 testing pairs, where some samples are under inconsistent illumination or color conditions. All of the images are available in three different resolutions. We select the full-resolution and half-resolution ones of training pairs to evaluate cross-domain generalization performance on large disparity. The evaluation metric is bad 2.0, where the percentage of the pixels with EPE larger than 2 pixels. ETH3D [57] is a gray-scale dataset with 27 training pairs and 20 testing pairs for indoor and outdoor scenes.
@@ IV-B Implemention Details
TABLE III: Quantitative comparisons on the KITTI 2012 [12] and KITTI 2015 [13] test set. We split the state-of-the-art methods into two parts according to the run-time. Our IGEV++ runs 16 update iterations at inference, and RT-IGEV++ runs 6 update iterations at inference. Bold: Best.
Target Method KITTI 2012 [12] KITTI 2015 [13] Run-time (ms)
@@
Real-Time HITNet [17] 2.00 2.65 1.41 1.89 0.4 0.5 1.74 3.20 1.98 20
@@
RT-IGEV++ (Ours) 1.93 2.51 1.29 1.68 0.4 0.5 1.48 3.37 1.79 48
Accuracy PSMNet [1] 2.44 3.01 1.49 1.89 0.5 0.6 1.86 4.62 2.32 410
@@
RAFT-Stereo [3] 1.92 2.42 1.30 1.66 0.4 0.5 1.58 3.05 1.82 380
GMStereo [5] - - - - - - 1.49 3.14 1.77 170
Croco-Stereo [6] - - - - - - 1.38 2.65 1.59 930
@@ IV-D Ablation Study
Effectiveness of MGEV. We explore the best settings for the MGEV and examine its effectiveness. For all models in these experiments, we perform 32 iterations of ConvGRUs update at inference. We take RAFT-Stereo [3] as our baseline, and replace its original backbone with MobileNetV2 100 for efficiency. As shown in Table II, the proposed GEV significantly improves prediction accuracy within the 192px disparity range (Disp<192px). Furthermore, the proposed MGEV can significantly enhance performance in large disparity regions by a substantial margin. The all-pairs correlations of RAFT-Stereo [3] (Baseline) lack non-local geometry and context information, and thus are difficult to handle local ambiguities in ill-posed regions. In contrast, our MGEV provides more comprehensive yet concise geometry information to ConvGRUs, yielding more effective optimization in each iteration. For each iteration, we index the fused geometry features from three ranges and three granularities of geometry encoding volumes. Our proposed selective geometry feature fusion module enables the model to better handle disparities of various ranges. The final model, referred to as IGEV++, achieves the best performance across all disparity ranges.
@@
Number of Iterations. Our IGEV++ achieves excellent performance even when the number of iterations is reduced. As shown in Table IV, we report the EPE (Disp<768px) of RAFT-Stereo [3], DLNR [16] and our IGEV++ on the Scene Flow test set [14]. Compared with all-pairs correlations in RAFT-Stereo and DLNR, our MGEV can facilitate more comprehensive yet concise geometry information to ConvGRUs, yielding more effective optimization in each iteration and thus effectively addressing local ambiguities in ill-posed regions. Thus when the number of iterations is reduced to 1, 2, 3 or 4, our IGEV++ outperforms RAFT-Stereo [3] and DLNR [16] with the same number of iterations by a large margin, e.g., surpassing RAFT-Stereo and DLNR by 72.91% and 57.27%, respectively, when the iteration number is 1. From Table IV and Figure 5, we can observe that our IGEV++ achieves state-of-the-art performance even with few iterations, providing users a flexible trade-off between time efficiency and accuracy according to their needs.
@@ IV-E Benchmark Results
KITTI. The comparison results with previous methods are shown in Table III. Our IGEV++ achieves the best performance among the published methods for almost all metrics on KITTI 2012 [12] and 2015 [13]. We first fine-tune the pre-trained Scene Flow [14] model on the virtual KITTI 2 [62] for 60k steps with a batch size of 8. Then we fine-tune it on the mixed dataset of KITTI 2012 [12] and KITTI 2015 [13] for 50k steps. The quantitative comparison results are shown in Table III. On KITTI 2012, our IGEV++ outperforms LEAStereo [61] and CREStereo [15] by 17.89% and 9.30% on 2-noc metric, respectively. On KITTI 2015, our IGEV++ surpasses GMStereo [5] and Croco-Stereo [6] by 14.69% and 9.58% on D1-all metric, respectively.
TABLE IV: Quantitative results on the Scene Flow test set for different iterations. The evaluation metric is EPE (Disp<768px). Our IGEV++ achieves better results than RAFT-Stereo with only 2 iterations compared to its 32 iterations.
@@
TABLE V: Evaluation in the reflective regions (ill-posed regions) of KITTI 2012 benchmark.
@@
TABLE VI: Quantitative comparisons with previous methods on ETH3D [57] and Middlebury [18] benchmarks. Bold: Best, Underline: Second best.
@@
Middlebury. Following CREStereo [15] and GMStereo [5], we first fine-tune the pre-trained Scene Flow model on the mixed Tartan Air [70], CREStereo dataset [15], Scene Flow [14], Falling Things [71], InStereo2k [72], CARLA HR-VS [73] and Middlebury [18] datasets using a crop size of $384\times 512$ for 200k steps. Then we fine-tune it on the mixed CREStereo dataset [15], Falling Things [71], InStereo2k [72], CARLA HR-VS [73] and Middlebury [18] datasets using a crop size of $384\times 768$ with a batch size of 8 for another 100k steps. As shown in Table. VI, our IGEV++ outperforms most published methods on Middlebury [18] benchmark. Visual comparisons are shown in Figure 7.
ETH3D. Following CREStereo [15] and GMStereo [5], we use a collection of several public stereo datasets for training. We first fine-tune the pre-trained Scene Flow model on the mixed Tartan Air [70], CREStereo dataset [15], Scene Flow [14], Sintel Stereo [74], InStereo2k [72] and ETH3D [75] datasets for 300k steps, and the crop size is $384\times 512$. Then we fine-tune it on the mixed CREStereo dataset [15], InStereo2k [72] and ETH3D [57] datasets for another 100k steps. As shown in Table VI, our IGEV++ outperforms most published methods on ETH3D [57] benchmark. Visual comparisons are shown in Figure 8, our IGEV++ performs better than RAFT-Stereo [3] and GMStereo [5] in textureless regions.
TABLE VII: Comparisons with other real-time methods on the Scene Flow test set [14]
@@
RT-IGEV++ (Ours, Iters=2) 0.60 36
RT-IGEV++ (Ours, Iters=4) 0.55 42
RT-IGEV++ (Ours, Iters=6) 0.52 48
RT-IGEV++ (Ours, Iters=8) 0.50 54
@@
TABLE VIII: Zero-shot generalization results on the large disparity dataset Middlebury. All models are trained on the synthetic Scene Flow dataset and then tested directly on the real Middlebury dataset. The evaluation metric is the 2-pixel outlier rate (Bad 2.0).
@@ IV-F Performance in Ill-posed Regions
To verify the performance of our IGEV++ in ill-posed regions, we compare our method with previously published state-of-the-art methods on reflective regions in the KITTI 2012 benchmark. As shown in Table V, our IGEV++ achieves the best performance among all methods. Compared to the all-pairs correlations of RAFT-Stereo, our MGEV encodes sufficient non-local geometry information and scene prior knowledge, which are essential for matching in ill-posed regions. Thus, our IGEV++ outperforms RAFT-Stereo [3] by 31.30% on the KITTI 2012 reflective regions (Table V) on the 3-noc metric. Visual comparisons are shown in Figure 9 that our IGEV++ performs very well in the reflective regions.

### v2

@@ III-C ConvGRU-based Update Operator
The objective of the update operator is to iteratively refine the initial disparity map using the indexed geometry features from MGEV. Before starting the iteration, we first apply $soft\;argmin$ to $\mathbf{G}^{s}$, $\mathbf{G}^{m}$, and $\mathbf{G}^{l}$ to regress a set of initial disparity maps $\mathbf{d}_{0}^{s}$, $\mathbf{d}_{0}^{m}$, and $\mathbf{d}_{0}^{l}$,
$\mathbf{d}_{0}^{s}=$ $\;\sum\limits_{d^{s}\in\mathcal{D}^{s}}d^{s}\times Softmax(\mathbf{G}^{s}(d^{s})),$ (5)
$\mathbf{d}_{0}^{m}=$ $\;\sum\limits_{d^{m}\in\mathcal{D}^{m}}d^{m}\times Softmax(\mathbf{G}^{m}(d^{m})),$
$\mathbf{d}_{0}^{l}=$ $\;\sum\limits_{d^{l}\in\mathcal{D}^{l}}d^{l}\times Softmax(\mathbf{G}^{l}(d^{l})).$
To speed up convergence, we take the initial disparity map $\mathbf{d}^{s}_{0}$ as the starting point for ConvGRU-based iterations. In each iteration, we utilize the previous disparity map to index a set of geometry features $\mathbf{f}_{G}^{s}$, $\mathbf{f}_{G}^{m}$, and $\mathbf{f}_{G}^{l}$ from $\mathbf{G}^{s}$, $\mathbf{G}^{m}$, and $\mathbf{G}^{l}$. We further propose a Selective Geometry Feature Fusion module (SGFF) to effectively integrate geometry features from different ranges and granularities of MGEV.
@@
$\mathbf{s}_{j}=$ $\;\sigma(\text{Conv}(\text{Concat}\{\mathbf{f}_{l,4},\mathbf{f}_{d}\})),j=s,m,l.$
@@
In each iteration, the fused geometry feature $\mathbf{f}_{G}$ and the previous disparity map $\mathbf{d}_{k-1}$ are passed through two encoder layers and then concatenated with $\mathbf{d}_{k-1}$ to form $x_{k}$. Then we use ConvGRUs to update the hidden state $h_{k-1}$ as RAFT-Stereo [3],
@@
where $W_{z}$, $W_{r}$, and $W_{h}$ are the parameters of the network, $c_{z}$, $c_{r}$, $c_{h}$ are context features generated from the context network. The number of channels for both the hidden states and the context features is 128. The $\text{Encoder}_{g}$ and $\text{Encoder}_{d}$ consist of two convolutional layers respectively. Based on the hidden state $h_{k}$, we decode a residual disparity $\triangle\mathbf{d}_{k}$ through two convolutional layers, then we update the disparity map,
@@
Figure 6: Visual comparisons with RAFT-Stereo [3] using few iterations for inference, with iteration numbers of 1, 2, and 3.
@@ III-E Loss Function
TABLE III: Quantitative comparisons between the predicted disparity maps for different ranges of geometry encoding volumes (GEV). The iterative small-range GEV performs well for small disparities but struggles with large disparities. On the other hand, the iterative large-range GEV handles large disparities effectively.
Model small-range medium-range large-range Scene Flow EPE
GEV GEV GEV Disp<192px Disp<384px Disp<768px
small-range GEV 0.47 0.87 1.06
medium-range GEV 0.51 0.67 0.82
large-range GEV 0.54 0.69 0.76
multi-range GEV (final) 0.43 0.56 0.67
Figure 8: Qualitative comparisons between the predicted disparity maps for different ranges of geometry encoding volumes (GEV). From top to bottom, the three inputs contain different maximum disparities: 662, 375, and 108, respectively. The second, third, and fourth columns show the predicted disparity maps and error maps generated by iterating over the small-range GEV, medium-range GEV, and large-range GEV, respectively. The final column shows the results of our final MGEV.
@@ IV-A Datasets and Evaluation Metrics
Middlebury V3 [18] is an indoor dataset, which provides 15 training pairs and 15 testing pairs, where some samples are under inconsistent illumination or color conditions. All of the images are available in three different resolutions. We select the full-resolution and half-resolution ones of training pairs to evaluate cross-domain generalization performance on large disparity. The evaluation metric is bad 2.0, where the percentage of the pixels with EPE is larger than 2 pixels. ETH3D [57] is a gray-scale dataset with 27 training pairs and 20 testing pairs for indoor and outdoor scenes.
@@ IV-B Implemention Details
TABLE IV: Quantitative comparisons on the KITTI 2012 [12] and KITTI 2015 [13] test sets. We categorize state-of-the-art methods into two groups based on their runtime and further into cost volume filtering-based (Filter.), transformer-based (Trans.), and iterative optimization-based (Iter.) methods. indicates that we used DepthAnythingV2-Large [59] was used as the feature extraction backbone, with frozen parameters during training. IGEV++ and IGEV++ run 16 update iterations at inference, and RT-IGEV runs 6 update iterations at inference. Bold: Best.
Target Category Method KITTI 2012 [12] KITTI 2015 [13] Run-time (ms)
@@
Real-Time Filter. HITNet [17] 2.00 2.65 1.41 1.89 0.4 0.5 1.74 3.20 1.98 20
@@
Iter. RT-IGEV (Ours) 1.93 2.51 1.29 1.68 0.4 0.5 1.48 3.37 1.79 48
Accuracy Filter. PSMNet [1] 2.44 3.01 1.49 1.89 0.5 0.6 1.86 4.62 2.32 410
@@
GANet+ADL [63] 1.52 2.01 0.98 1.29 0.4 0.5 1.38 2.38 1.55 670
Trans. GMStereo [5] - - - - - - 1.49 3.14 1.77 170
Croco-Stereo [6] - - - - - - 1.38 2.65 1.59 930
Iter. RAFT-Stereo [3] 1.92 2.42 1.30 1.66 0.4 0.5 1.58 3.05 1.82 380
DLNR [16] - - - - - - 1.60 2.59 1.76 -
@@
Selective-IGEV [64] 1.59 2.05 1.07 1.38 0.4 0.4 1.33 2.61 1.55 240
TC-Stereo [65] - - - - - - 1.29 2.33 1.46 90
ViTAStereo [66] 1.46 1.80 0.93 1.16 0.4 0.4 1.21 2.99 1.50 360
IGEV (Ours) 1.71 2.17 1.12 1.44 0.4 0.4 1.38 2.67 1.59 180
@@ IV-D Ablation Study
Effectiveness of MGEV. We explore the best settings for the MGEV and examine its effectiveness. For all models in these experiments, we perform 32 iterations of ConvGRUs update at inference. We take RAFT-Stereo [3] as our baseline, and replace its original backbone with MobileNetV2 100 for efficiency. As shown in Table II, the proposed GEV significantly improves prediction accuracy within the 192px disparity range (Disp<192px). Furthermore, the proposed MGEV can significantly enhance performance in large disparity regions by a substantial margin. The all-pairs correlations of RAFT-Stereo [3] (Baseline) lack non-local geometry and context information, and thus make it difficult to handle local ambiguities in ill-posed regions. In contrast, our MGEV provides more comprehensive yet concise geometry information to ConvGRUs, yielding more effective optimization in each iteration. For each iteration, we index the fused geometry features from three ranges and three granularities of geometry encoding volumes. Our proposed selective geometry feature fusion module enables the model to better handle disparities of various ranges. The final model, referred to as IGEV++, achieves the best performance across all disparity ranges.
Analysis of GEV at different ranges. To validate the impact of different ranges of GEV on the final disparity results, we separately constructed small-range GEV, medium-range GEV, and large-range GEV, with the iterative optimization operations consistent with those in IGEV++. Accordingly, during the iterative process, the geometry feature $\mathbf{f}_{G}$ is derived from the small-range GEV, medium-range GEV, and large-range GEV, respectively. As can be seen from Table III and Figure 8, the iterative small-range GEV performs well for small disparities but struggles with large disparities. On the other hand, the iterative large-range GEV handles large disparities effectively; however, due to coarse-grained matching, it tends to lose details in small-range disparity regions. Our MGEV combines the strengths of both, achieving the best performance.
Adaptive Patch Matching. When adaptive patch matching (APM) is not used, MGEV involves constructing multi-size cost volumes at 1/4 resolution. Specifically, the large-size cost volume contains 192 disparity candidates (768/4) for cost aggregation, the medium-size cost volume contains 96 disparity candidates (384/4), and the small-size cost volume contains 48 disparity candidates (192/4) for cost aggregation. The use of APM enables the efficient construction of MGEV, where only 48 disparity candidates (768/4/4) are used for cost aggregation in the large-range GEV, as well as in the medium disparity range GEV. Moreover, a large-size (192 disparity candidates) cost volume significantly increases the complexity of cost aggregation and disparity regression, particularly in extensive textureless regions and areas with large disparities, as shown in Figure 10 and Table II.
@@
Number of Iterations. Our IGEV++ achieves excellent performance even when the number of iterations is reduced. As shown in Table V, we report the EPE (Disp<768px) of RAFT-Stereo [3], DLNR [16] and our IGEV++ on the Scene Flow test set [14]. Compared with all-pairs correlations in RAFT-Stereo and DLNR, our MGEV can facilitate more comprehensive yet concise geometry information to ConvGRUs, yielding more effective optimization in each iteration and thus effectively addressing local ambiguities in ill-posed regions. Thus when the number of iterations is reduced to 1, 2, 3 or 4, our IGEV++ outperforms RAFT-Stereo [3] and DLNR [16] with the same number of iterations by a large margin, e.g., surpassing RAFT-Stereo and DLNR by 72.91% and 57.27%, respectively, when the iteration number is 1. From Table V and Figure 5, we can observe that our IGEV++ achieves state-of-the-art performance even with few iterations, providing users a flexible trade-off between time efficiency and accuracy according to their needs.
Figure 10: Qualitative comparisons with/without APM module. Row 1 shows performance on the Scene Flow test set, while Row 2 presents the zero-shot generalization performance on the Middlebury dataset. Our APM is more robust in handling large ill-posed regions and large disparities.
@@ IV-E Benchmark Results
KITTI. The comparison results with previous methods are shown in Table IV. Our IGEV++ achieves the best performance among the published methods for almost all metrics on KITTI 2012 [12] and 2015 [13]. We first fine-tune the pre-trained Scene Flow [14] model on the virtual KITTI 2 [67] for 60k steps with a batch size of 8. Then we fine-tune it on the mixed dataset of KITTI 2012 [12] and KITTI 2015 [13] for 50k steps. The quantitative comparison results are shown in Table IV. On KITTI 2012, our IGEV++ outperforms LEAStereo [62] and CREStereo [15] by 17.89% and 9.30% on 2-noc metric, respectively. On KITTI 2015, our IGEV++ surpasses GMStereo [5] and Croco-Stereo [6] by 14.69% and 9.58% on D1-all metric, respectively. As shown in Table IV, IGEV++ increases the runtime by 100ms compared to IGEV. Prior to iterative optimization, IGEV++ incurs an additional 40ms for constructing and filtering the medium-range and large-range cost volumes, whereas IGEV only constructs and filters a single small-range cost volume. During the iterative process (16 iterations during inference), IGEV++ further introduces overhead by indexing a set of geometry features $\mathbf{f}_{G}^{m}$ and $\mathbf{f}_{G}^{l}$ from the medium-range GEV and large-range GEV at each iteration, and by computing selective weights to fuse these geometry features. This step contributes an additional 60ms to the overall runtime.
TABLE V: Quantitative results on the Scene Flow test set for different iterations. The evaluation metric is EPE (Disp<768px). Our IGEV++ achieves better results than RAFT-Stereo with only 2 iterations compared to its 32 iterations.
@@
TABLE VI: Evaluation in the reflective regions (ill-posed regions) of KITTI 2012 benchmark.
@@
TABLE VII: Quantitative comparisons with previous methods on ETH3D [57] and Middlebury [18] benchmarks. Bold: Best, Underline: Second best.
@@
Middlebury. Following CREStereo [15] and GMStereo [5], we first fine-tune the pre-trained Scene Flow model on the mixed Tartan Air [74], CREStereo dataset [15], Scene Flow [14], Falling Things [75], InStereo2k [76], CARLA HR-VS [77] and Middlebury [18] datasets using a crop size of $384\times 512$ for 200k steps. Then we fine-tune it on the mixed CREStereo dataset [15], Falling Things [75], InStereo2k [76], CARLA HR-VS [77] and Middlebury [18] datasets using a crop size of $384\times 768$ with a batch size of 8 for another 100k steps. As shown in Table. VII, our IGEV++ outperforms most published methods on Middlebury [18] benchmark. Visual comparisons are shown in Figure 7.
ETH3D. Following CREStereo [15] and GMStereo [5], we use a collection of several public stereo datasets for training. We first fine-tune the pre-trained Scene Flow model on the mixed Tartan Air [74], CREStereo dataset [15], Scene Flow [14], Sintel Stereo [78], InStereo2k [76] and ETH3D [79] datasets for 300k steps, and the crop size is $384\times 512$. Then we fine-tune it on the mixed CREStereo dataset [15], InStereo2k [76] and ETH3D [57] datasets for another 100k steps. As shown in Table VII, our IGEV++ outperforms most published methods on ETH3D [57] benchmark. Visual comparisons are shown in Figure 9, our IGEV++ performs better than RAFT-Stereo [3] and GMStereo [5] in textureless regions.
TABLE VIII: Comparisons with other real-time methods on the Scene Flow test set [14]
@@
RT-IGEV (Ours, Iters=2) 0.60 36
RT-IGEV (Ours, Iters=4) 0.55 42
RT-IGEV (Ours, Iters=6) 0.52 48
RT-IGEV (Ours, Iters=8) 0.50 54
@@
TABLE IX: Zero-shot generalization results on the large disparity dataset Middlebury V3. All models are trained on the synthetic Scene Flow dataset and then tested directly on the real Middlebury V3. The evaluation metric is the 2-pixel outlier rate (Bad 2.0).
@@ IV-F Performance in Ill-posed Regions
To verify the performance of our IGEV++ in ill-posed regions, we compare our method with previously published state-of-the-art methods on reflective regions in the KITTI 2012 benchmark. As shown in Table VI, our IGEV++ achieves the best performance among all methods. Compared to the all-pairs correlations of RAFT-Stereo, our MGEV encodes sufficient non-local geometry information and scene prior knowledge, which are essential for matching in ill-posed regions. Thus, our IGEV++ outperforms RAFT-Stereo [3] by 31.30% on the KITTI 2012 reflective regions (Table VI) on the 3-noc metric. Visual comparisons are shown in Figure 11 that our IGEV++ performs very well in the reflective regions.